### v6

# Advancing Prompt Learning through an External Layer

Abstract.
Prompt learning represents a promising method for adapting pre-trained vision-language models (VLMs) to various downstream tasks by learning a set of text embeddings. One challenge inherent to these methods is the poor generalization performance due to the invalidity of the learned text embeddings for unseen tasks. A straightforward approach to bridge this gap is to freeze the text embeddings in prompts, which results in a lack of capacity to adapt VLMs for downstream tasks. To address this dilemma, we propose a paradigm called EnPrompt with a novel External Layer (EnLa). Specifically, we propose a textual external layer and learnable visual embeddings for adapting VLMs to downstream tasks. The learnable external layer is built upon valid embeddings of pre-trained CLIP. This design considers the balance of learning capabilities between the two branches. To align the textual and visual features, we propose a novel two-pronged approach: i) we introduce the optimal transport as the discrepancy metric to align the vision and text modalities, and ii) we introduce a novel strengthening feature to enhance the interaction between these two modalities. Four representative experiments (i.e., base-to-novel generalization, few-shot learning, cross-dataset generalization, domain shifts generalization) across 15 datasets demonstrate that our method outperforms the existing prompt learning method.
Vision-language Model, Prompt Learning, Optimal Transport
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680953isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
1. Introduction
Vision-language models (VLMs), such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021), have demonstrated remarkable generalization performance for various downstream tasks (Xie et al., 2024, 2023, 2022). VLMs are typically trained to align textual and visual modalities using large-scale datasets, which allows them to encode open-vocabulary (Fang et al., 2021, 2020) concepts in a shared embedding space. Thanks to the modality matching ability, CLIP has achieved remarkable success across various downstream tasks (Greer et al., 2024; Luo et al., 2018, 2023c; Zhou et al., 2022a; Luo et al., 2023d; Guo et al., 2024), such as action recognition (Wang et al., [n. d.]), image generation (Sanghi et al., 2021), and image segmentation (Luo et al., 2023a; Zheng Ding, 2023). One of the attractive features of CLIP is the ability to perform zero-shot inference, where some pre-defined text inputs (a.k.a. prompts) are used to generate classification weights for predicting image features during inference.
Seminal explorations involve hand-crafted text prompts, e.g., “a photo of a [class]” as the prompt for text encoder. Advanced works introduce a set of learnable parameters to adapt VLMs to downstream tasks. For instance, CoOp (Zhou et al., 2022c) keeps the weights of CLIP frozen while learning the text embeddings of prompts for efficient task-specific adaptation. These prompt-learning approaches achieve significant performance improvements over manually tuned prompts on various typical scenarios (Chen et al., 2023; Cao et al., 2024c; Etchegaray et al., 2024). However, the learned text embeddings typically produce worse performance compared with hand-crafted prompts (Radford et al., 2021) for unseen tasks. Specifically, applying these learned text prompts to unseen classes leads to degenerated model generalization performance. Recent works mitigate this problem through the overfitting view. To overcome the challenge, an image-conditional prompt (CoCoOp) learning approach (Zhou et al., 2022b) is proposed to improve the generalization performance on unseen classes. Unfortunately, CoCoOp exhibits lower generalization performance than hand-crafted zero-shot CLIP when applied to novel classes, as depicted in Table 4. This could be attributed to the fact that learned embeddings are usually invalid on novel classes.
To address this dilemma, we propose a paradigm called EnPrompt with a novel External Layer (EnLa). Specifically, we propose adding a textual external layer on top of frozen text embeddings and using learnable visual embeddings for adapting VLMs to downstream tasks. To align the textual and visual features, we introduce the optimal transport (Villani et al., 2009) as a discrepancy metric that measures the difference between the visual and textual features, where the optimal transport can calculate the distance between two distributions under the form of multiple sampling. Further, we connect the text and image encoders through a strengthening feature instead of ordinary coupling for the dimension transformation of learnable hidden vectors. As shown in Figure 1, EnPrompt outperforms the existing state-of-the-art method on the base-to-novel generalization task. Our main contributions can be summarized as follows:
  • We address the performance-degeneration issue of VLMs on unseen tasks by introducing a novel External Layer (EnLa) and freezing the text embeddings.
  • We align the visual and textual features of the two modalities by introducing a novel two-pronged approach. Specifically, we introduce the optimal transport as the discrepancy metric to measure and mitigate the difference between the visual and textual features. Meanwhile, we introduce a novel interaction strategy between modalities to strengthen the modality fusion.
  • Four highly representative experiments demonstrate that our method can consistently outperform the existing methods across 15 datasets, achieving state-of-the-art performance.
2. Methodology
Figure 1. Performance comparison on base-to-novel generalization. EnPrompt (Ours) outperforms previous state-of-the-art methods on 11 datasets.
2.1. Preliminaries
We provide a brief introduction to vision-language pre-training, with a specific focus on CLIP, which is a zero-shot learning approach without fine-tuning. We build our approach based on CLIP. CLIP has a text encoder $\mathcal{F}_{t}(\cdot)$ and an image encoder $\mathcal{F}_{v}(\cdot)$, which separately map a textual input (a.k.a. prompt) $\mathbf{p}$ and a visual input, i.e., an image, $\boldsymbol{x}$ into a shared feature space through many transformer blocks. The outputs of two encoders are denoted as ${\boldsymbol{t}}=\mathcal{F}_{t}(\mathbf{p})$ and ${\boldsymbol{v}}=\mathcal{F}_{v}(\mathbf{x})$. The image encoder aims to transform the input images into feature embeddings, while the text encoder generates representations for word embedding sequences.
Figure 2. Comparison of our method with CoOp and MaPLe. Our method freezes the textual embeddings and learns an External Layer (EnLa) on top of it. Optimal transport is introduced to align visual and textual modalities apart from the conventional cross entropy (CE) loss. The textual and visual encoders are connected through a strengthening feature. This feature and the learnable visual prompt are fused and utilized for deep prompting of the visual encoder.
During the pre-training phase of CLIP, these two encoders are simultaneously trained on large-scale datasets of text-image pairs, where a contrastive loss is employed to maximize the cosine similarity of text-image pairs and minimize the cosine similarity between unmatched pairs in the feature space. The final prediction probability of alignment is computed by the matching score as follows:
(1) $p(y\mid\boldsymbol{x},\mathcal{P})=\frac{\exp\left\{sim\left(\mathcal{F}_{t}\left(\mathbf{p}_{y}\right),\mathcal{F}_{v}(\mathbf{x})\right)/\tau\right\}}{\sum_{\mathbf{p}_{i}\in\mathcal{P}}\exp\left\{sim\left(\mathcal{F}_{t}\left(\mathbf{p}_{i}\right),\mathcal{F}_{v}(\mathbf{x})\right)/\tau\right\}},$
where $y\in\mathcal{Y}$ is the label of $\mathbf{x}\in\mathcal{X}$, $\mathcal{P}=\left\{\mathbf{p}_{i}\right\}_{i=1}^{C}$ denotes the set of $C$ pre-defined prompts, $sim(\cdot,\cdot)$ stands for cosine similarity between two vectors, and $\tau>0$ represents a temperature parameter. Here, the classifier consists of $C$ textual features derived from pre-defined prompts $\mathcal{P}=\left\{\mathbf{p}_{i}\right\}_{i=1}^{C}$, where the prompt $\mathbf{p}_{i}$ for the $i$-th class may have the form of “a photo of a [class]”.
Advanced works take a further step to investigate the possibility of aligning images and prompts. The insight of these works is that the prompts tuning for given images could be superior to hand-crafted prompts. Specifically, the class name is retained as prior knowledge to ensure the learned prompts can form a classifier, while the word (a.k.a. context) embeddings of prompts are modeled as learnable parameters. Here, the learnable words in the above prompt are typically initialized using “a photo of a []”. The embeddings optimization approach can be formalized as follows,
(2) $\min_{\mathbf{w}}\ell(\mathbf{w})=-\log p(y\mid\mathbf{x},\mathcal{P}(\mathbf{w})),$
where $\mathbf{w}$ stands for learnable embeddings used to model the context in prompts. The prompt $\mathbf{p}_{i}\in\mathcal{P}(\mathbf{w})$ learned by the context optimization approach may have the form of,
[…]
where ⓒ is the concatenating operation. Following CoOp (Zhou et al., 2022c), the embeddings $\mathbf{w}$ can be shared across classes. In the inference phase, the prompts with the learned embeddings can produce textual features for classification.
2.2. EnLa with Frozen Text Embeddings
As depicted in Table 4, the existing methods CoOp (Zhou et al., 2022c) and CoCoOp (Zhou et al., 2022b) work well on the base classes observed during training and beat the hand-crafted prompts method employed by CLIP (Radford et al., 2021) by a significant margin. However, these methods typically perform worse than hand-crafted prompts for novel classes (unseen tasks). This may be attributed to the non-adaptive of the learned text embeddings on unseen tasks. Learning task-specific text embeddings can result in the loss of the essential general textual knowledge having a strong generalization ability.
Frozen Text Embeddings. Built upon the observation, we propose to freeze the text embeddings of prompts for unseen tasks. Specifically, the text embeddings are set to be frozen and non-learnable, which aims to release the potential generalization ability of CLIP and ensure pre-trained ability for unseen tasks. Here, the input words in the prompt are typically initialized using “a photo of a”, which can be vectorized into embeddings. In our upcoming ablation experiments, we plan to incorporate other manual template inputs for validation. However, freezing text embeddings of prompts will cause limited model capacity, which results in a lack of capacity to adapt VLMs for downstream tasks.
External Layer (EnLa). To address this dilemma, we propose to introduce an External Layer (EnLa) of the text branch and learnable visual embeddings of the visual branch to learn with VLMs for adapting downstream tasks. This design considers the balance of learning capabilities between the two branches. As depicted in Figure 2, EnLa can be considered as an extension of the text encoder of VLMs, which is an auxiliary layer introduced to the architecture. Introducing the EnLa will make VLMs learnable while keeping the text embeddings frozen. In contrast to CLIP (Radford et al., 2021), our method is being explored in the field of few-shot learning.
Let $\mathcal{E}_{\boldsymbol{\theta}}(\cdot)$ denote the EnLa parameterized by $\boldsymbol{\theta}$, which is updated for adapting VLMs to downstream tasks. Moreover, let $\mathcal{P}$ denote the frozen text embeddings. The $\mathcal{E}_{\boldsymbol{\theta}}(\cdot)$ is also realized as an incomplete embedding transformer, showing the same spirit with neural networks. Specifically, let $\boldsymbol{e}$ denote the output of the EnLa,
(4) $\boldsymbol{e}=\mathcal{E}_{\boldsymbol{\theta}}(\mathcal{P}),$
where $\boldsymbol{e}$ is further transferred as input to two directions: text encoder $\mathcal{F}_{t}(\cdot)$ and vision-language interaction, particularly towards the image encoder $\mathcal{F}_{v}(\cdot)$.
EnLa Design Analysis. We compare the performance of a single layer with a two-layer, referring to the bottleneck structure (Linear-ReLU-Linear), on 11 datasets with 16 shots for base-to-novel tasks. HM refers to the harmonic mean (Xian et al., 2017). The results in Table 1 indicate that the single 512x512 structure (row-5) provides better performance. It achieves the highest HM of $80.64\%$, which surpasses all reduction factors of the hidden layers that have 32x, 16x, 8x, and 4x. Thus, we used a single layer of EnLa in all of our experiments.
Table 1. Ablation study on the different designs of EnLa. Results are averaged over 11 datasets.
Layer Design Reduction factor Base Acc Novel Acc. HM
1: (512 x 16) (16 x 512) 32x 83.9 73.40 78.3
2: (512 x 32) (32 x 512) 16x 84.6 76.8 80.51
3: (512 x 64) (64 x 512) 8x 84.45 76.70 80.41
4: (512 x 128) (128 x 512) 4x 84.40 77.0 80.54
5: (512 x 512) / 84.71 76.90 80.64
2.3. Alignment with Optimal Transport
To align two modalities for generalization, we introduce the optimal transport as a discrepancy metric and formulate the feature sets as discrete probability distributions for generalization performance. In contrast to conventional distance metrics such as Euclidean distance, optimal transport (Villani et al., 2009) learns an adaptive transport plan to calculate the cross-modality distance, facilitating fine-grained matching across the two modalities. It enables us to align visual features in a more precise and adaptive manner, capturing subtle nuances and enhancing the matching accuracy between modalities.
Optimal Transport. The Optimal Transport (OT) distance is a commonly employed metric for comparing distributions. In the context of our framework, we specifically concentrate on the discrete situation as it aligns more closely with our approach. In this scenario, we consider two sets of points or features.
Given two sets that contain $N$ and $M$ points respectively, the discrete distributions can be formulated as follows:
(5) $\mathbf{Z}_{\theta}=\sum_{n=1}^{N}\theta_{n}\delta_{\boldsymbol{e}_{n}}\quad\text{ and }\quad\mathbf{Q}_{\beta}=\sum_{m=1}^{M}\beta_{m}\delta_{\boldsymbol{l}_{m}}$
where $\boldsymbol{\theta}\in\Delta^{N}$ and $\boldsymbol{\beta}\in\Delta^{M}$ are discrete probability vectors that sum to 1 , and $\delta_{e}$ refers to a point mass located at point $e$ in the embedding space. The OT distance between $\mathbf{Z}_{\theta}$ and $\mathbf{Q}_{\beta}$ is defined as:
(6) $d_{\mathbf{OT}}(\boldsymbol{\theta},\boldsymbol{\beta}):=\min_{\mathbf{T}}<\mathbf{T},\mathbf{C}>,$
[…]
where <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the Frobenius dot-product and $\mathbf{1}_{N}$ is the $N$ dimensional vector of ones. $\mathbf{C}\in\mathbb{R}_{>0}^{N\times M}$ is the cost matrix of the transport, and $C_{nm}$ denotes the transport cost between points $\boldsymbol{e}_{n}$ and $\boldsymbol{l}_{m}$, such as the cosine distance $C_{nm}=1-\operatorname{cos}\left(\boldsymbol{e}_{n},\boldsymbol{l}_{m}\right).\mathbf{T}\in\mathbb{R}_{>0}^{N\times M}$ denotes the transport plan to be learned. OT distance is then minimized over all the joint probabilities of $N\times M$ space with two marginal constraints. As computing the above OT distance has the cubic time complexity, we apply the Sinkhorn distance  (Cuturi, 2013) that regularizes with an entropic constraint:
(8) $d_{\mathbf{OT},\lambda}(\boldsymbol{\theta},\boldsymbol{\beta})=$ $d_{\mathbf{OT}}(\boldsymbol{\theta},\boldsymbol{\beta})-\lambda h(\mathbf{T}),$
[…]
where $h(\mathbf{T})$ is the entropy of transport plan $\mathbf{T}$ and $\lambda\geq 0$ is a hyper-parameter. It can be optimized within a few iterations by the Sinkhorn algorithm with the Lagrange multiplier of the entropy constraint.
Modalities Alignment with Optimal Transport. The transport plan is efficiently computed through a limited number of matrix multiplications as a forward module. These matrix multiplications are crucial for determining the gradients that are then preserved for back-propagation.
Specifically, let $\boldsymbol{v}$ denote the visual feature and $\boldsymbol{t}$ denote the textual feature. The output of the image encoder is a tensor with the shape, where $H$ and $W$ are the height and width. Therefore, we can obtain $M=H\times W$ local visual features. Further, let $\boldsymbol{V}=\left\{\left.\boldsymbol{v}_{m}\right|_{m=1}^{M}\right\}$ denote the local visual features as the fixed set, let $\boldsymbol{t}_{k}$ denote the text feature as the fixed set for class ${k}$. Our method learns the transport plan $\mathbf{T}$ by minimizing the following OT distance to push $\boldsymbol{t}_{k}$ to $\boldsymbol{V}$ for fine-grained alignment:
(10) $d_{\mathrm{OT}}(k)=d_{\mathrm{OT}}\left(\boldsymbol{\theta},\boldsymbol{\beta}\mid\mathbf{1}-\boldsymbol{V}^{\intercal}\boldsymbol{t}_{k}\right),$
where $\boldsymbol{C}=\mathbf{1}-\boldsymbol{V}^{\intercal}\boldsymbol{t}_{k}$ denotes that we use the cosine distance between $\boldsymbol{V}$ and $\boldsymbol{t}_{k}$ as the cost matrix. Then, we can obtain the solution of transport plan $\mathbf{T}^{*}$ and the final OT distance $d_{\mathrm{OT}}(k)$. Given the OT distance between $\boldsymbol{t}_{k}$ and $\boldsymbol{V}$, and image $\boldsymbol{x}$, we reformulate the final prediction probability of V-L alignment as follows:
(11) $p_{\mathrm{OT}}(y=k\mid\boldsymbol{x})=\frac{\exp\left(\left(1-d_{\mathrm{OT}}(k)\right)/\tau\right)}{\sum_{k^{\prime}=1}^{K}\exp\left(\left(1-d_{\mathrm{OT}}\left(k^{\prime}\right)\right)/\tau\right)}\text{. }$
In our method, we first fix the visual and textual features to optimize the optimal transport problem to calculate the cross-modality distance, obtaining the transport plan $\mathbf{T}^{*}$. Further, we back-propagate the gradient with cross-entropy loss to learn the learnable parameters of our method by fixing $\mathbf{T}^{*}$. This process is more robust to variations in visual domain shift tasks and tolerant to generalization.
2.4. Alignment with Strengthening Feature
In order to align two modalities for unseen tasks, we propose to connect the visual and textual modalities through a novel strengthening feature with a vision-language (V-L) function instead of coupling learnable hidden vectors for V-L alignment.
We introduce learnable visual embeddings for enhancing learnability of VLMs for downstream tasks. In contrast to VPT (Jia et al., 2022), our method employs a multi-modal prompt design. Specifically, let $\mathcal{T}_{\epsilon}(\cdot)$ denote the V-L function parameterized by $\epsilon$ to transfer features from textual branch to visual branch. The $\mathcal{T}_{\epsilon}(\cdot)$ is a function net with dimensionality of [512, 768] for dimension alignment of the visual branch and text branch. The input feature of the V-L function is the output produced by the EnLa. Further, we append $L$ learnable visual input embeddings $\mathcal{P}_{v}$, which is a vector with [$L$, 768], given as follows,
(12) $\mathcal{P}_{v}=\left\{\boldsymbol{p}_{\boldsymbol{v}}^{1},\boldsymbol{p}_{\boldsymbol{v}}^{2},\cdots,\boldsymbol{p}_{\boldsymbol{v}}^{L}\right\}.$
To analyze generalization capabilities of alignment, we are conducting an evaluation to assess the performance of different connection positions of the image encoder, as depicted in Table 2.
Table 2. Comparisons of different positions of strengthening feature in base-to-novel generalization. Results are averaged over 11 datasets.
Position of Strengthening Feature Base Acc. Novel Acc. HM
1: Deep layer of image encoder 83.85 76.31 79.90
2: Input layer of image encoder 84.71 76.90 80.64
The deep layer (row-1) of the image encoder is the hidden layer of the image encoder, without visual input (first) layer. The visual input layer (row-2) of the image encoder method is focused on the initialization of the visual embeddings. These designs combine prompting benefits in both branches by enforcing the image encoder representation ability through text knowledge transformation. The position (row-2) is more effective than the position (row-1) in unseen tasks. It can be attributed to that position (row-1) lacks visual embedding initialization, leading to a limited ability to adapt the image encoder with the unbalance of V-L alignment. As a consequence, in our subsequent comprehensive experiments, we adopted the connection position with the input layer of the image encoder (row-2) in our approach. Let $\boldsymbol{p}_{\boldsymbol{v}}(\boldsymbol{e})$ denote the fusion embeddings, it can be expressed as:
(13) $\boldsymbol{p}_{\boldsymbol{v}}(\boldsymbol{e})=\boldsymbol{p}_{\boldsymbol{v}}+\mathcal{T}_{\boldsymbol{\epsilon}}(\boldsymbol{e}),$
the $\boldsymbol{e}$ is the output of EnLa, and the fusion embeddings are further introduced in the transformer of the image encoder for deep prompting.
In contrast to EnPrompt (Ours), CoCoOp transfers the output features of the image encoder to the text embeddings, which is more inference time consumption. MaPLe initializes the multi-learnable hidden vectors for V-L alignment and couples the text embeddings to the visual encoder. The coupling function maps these initialized hidden vectors of deep layer and text embeddings with the image encoder and text encoder for more parameters. However, our approach introduces the learnable visual initialization embeddings of the image encoder with the strengthening feature fusion. However, MaPLe learns to visual branch with the coupled visual embeddings without the visual embeddings initialization step, which limits the ability of the visual branch to adapt downstream data distributions.
3. Experiments
3.1. Benchmark Setting
We compare the performance with CLIP (Radford et al., 2021), CoOp (Zhou et al., 2022c), Clip-Adapter (Gao et al., 2023), CoCoOp (Zhou et al., 2022b), PLOT (Chen et al., 2022), and MaPLe (Khattak et al., 2023). Please note that CLIP was originally a zero-shot method. However, in this context, we linearly process it using a few-shot method. The Clip-Adapter (Gao et al., 2023) is a frozen prompt method with the adapter component behind the encoder, which is different from EnPrompt. We compared it with PLOT (Chen et al., 2022), which is an optimal transport method. PLOT optimizes four sentences simultaneously (“a photo of a dog”, “a picture of a dog”, “a drawing of a dog”, “a good drawing of a dog”). The output text feature of EnPrompt is global feature, the input sentence is one text prompt such as “a photo of a”, and the text prompt is non-learnable.
Few-shot Learning. We employ this setup to evaluate EnPrompt in conditions of highly restricted supervision. We evaluate the model on 11 datasets by conducting tests with varying K-shots per class, where K takes the values of 1, 2, 4, 8, and 16.
Base-to-Novel Generalization. We evaluate the ability to generalize following the setting where the datasets are split into base and novel classes. The model is trained only on the base classes in 16 shots setting and evaluated on base and novel classes.
Cross-dataset Evaluation. To validate the potential of our approach in cross-dataset transfer, we evaluate our ImageNet-trained model in 16 shots directly on other 10 datasets. This experiment aims to verify whether our model can successfully complete the unseen generalization.
Domain Generalization. We evaluate the robustness of our approach to domain shift datasets. Similar to cross-dataset evaluation, we train our model using the ImageNet in 16 shots and evaluate its performance directly on 4 different variants of the ImageNet.
Datasets. For base to novel class generalization, few-shot settings, and cross-dataset evaluation, we use 11 image recognition datasets. The datasets cover multiple recognition tasks including ImageNet (Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) which consists of generic objects, OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flowers102 (Nilsback and Zisserman, 2008), Food101 (Bossard et al., 2014), and FGVCAircraft  (Maji et al., 2013) for fine-grained classification, SUN397 (Xiao et al., 2010) for scene recognition, UCF101 (Soomro et al., 2012) for action recognition, DTD  (Cimpoi et al., 2014) for texture classification, and EuroSAT  (Helber et al., 2019) which consists of satellite images. For domain generalization benchmark, we use ImageNetA (Hendrycks et al., 2021b), ImageNet-R  (Hendrycks et al., 2021a), ImageNet-Sketch (Wang et al., 2019) and ImageNetV2 (Recht et al., 2019) as domain shift datasets.
Implementation Details. We use a ViT-B/16-based CLIP model, and report results averaged over 3 runs. We set the learnable visual embedding length to 4. We use deep prompting (Khattak et al., 2023) for the image encoder. Training for 50 epochs for few-shot setting, 20 epochs for domain generalization setting and cross-dataset evaluation setting. In the base-to-novel setting, we apply 30 epochs. We use an SGD optimizer with a learning rate of 0.0025. All experiments are run on a single Nvidia A6000 GPU. For domain generalization and cross-dataset evaluation, we train the ImageNet source model on all classes with K = 16 shots in the first 3 transformer layers of the image encoder. For the few-shot learning and base-to-novel tasks, we set visual embeddings learning depth to 9. We initialize the text embeddings using the hand-crafted words of “a photo of a []”.
Figure 3. Few-shot Learning Experiments. All methods are trained on the ViT-B/16 CLIP backbone. EnPrompt (Ours) demonstrates consistent improvements over existing methods, specifically for minimal training data such as K=2. On average, EnPrompt provides the highest performance gains for all shots (K = 1,2,4,8,16).
3.2. Effect of Components
In Table 3, we disentangle the components and show the individual contributions to the base-to-novel generalization task. Results are averaged over 11 datasets with 16-shot. HM refers to harmonic mean. Integrating EnLa (row-2) outperforms baseline methods (row-1) on novel classes while maintaining base class gains. By enforcing visual embeddings (row-3), HM performance significantly increases due to the ability to adapt the image encoder for better V-L alignment. Integrating optimal transport (row-4) outperforms component (row-3) with an improvement of $1.08\%$ in HM. This suggests that the regularization of OT is more effective than the traditional metric function in calculating the cross-modality distance. Finally, combined with strengthening features and connecting to overcome the mismatch between the text and visual branch (row-5), our method achieves improvements on both base and novel classes.
Table 3. Effect of our components in base-to-novel generalization. Results are averaged over 11 datasets.
Components Base Acc. Novel Acc. HM
1: Frozen text embeddings 69.34 74.22 71.70
2: + EnLa 80.24 74.43 77.25
3: + Visual embedding learning 82.25 75.94 78.98
4: + Optimal transport 84.50 76.05 80.06
5: + Strengthening feature 84.71 76.90 80.64
Table 4. Base-to-novel generalization experiments. EnPrompt (Ours) demonstrates strong generalization results over existing methods on 11 recognition datasets. Here, the CLIP refers to the linear probe CLIP.
Dataset CLIP CoOp CLIP-Adapter CoCoOp PLOT MaPLe Ours $\Delta$
Average Base 69.34 82.69 82.91 80.47 81.3 82.28 84.71 +2.4
Novel 74.22 63.22 63.98 71.69 72.2 75.14 76.90 +1.8
HM 71.70 71.66 72.23 75.83 76.48 78.55 80.64 +2.1
ImageNet Base 72.43 76.47 76.88 75.98 75.33 76.66 77.70 +1.1
Novel 68.14 67.88 68.1 70.43 70.48 70.54 70.65 +0.1
HM 70.22 71.92 72.23 73.10 72.83 73.47 74.07 +0.6
Caltech101 Base 96.84 98.00 98.1 97.96 97.86 97.74 98.40 +0.7
Novel 94.00 89.81 90.00 93.81 93.99 94.36 94.07 -0.3
HM 95.40 93.73 93.89 95.84 95.92 96.02 96.2 +0.2
OxfordPets Base 91.17 93.67 93.88 95.20 95.7 95.43 95.67 +0.2
Novel 97.26 95.29 95.55 97.69 98.1 97.76 97.63 -0.1
HM 94.12 94.47 94.74 96.43 76.80 96.58 96.67 +0.1
StanfordCars Base 63.37 78.12 78.35 70.49 71.5 72.94 78.70 +5.8
Novel 74.89 60.40 60.55 73.59 73.77 74.00 75.67 +1.6
HM 68.65 68.13 68.33 72.01 72.62 73.47 77.22 +3.8
Flowers102 Base 72.08 97.60 97.61 94.87 95.1 95.92 98.47 +2.5
Novel 77.80 59.67 59.98 71.75 72.2 72.46 77.00 +4.4
HM 74.83 74.06 74.32 81.71 82.10 82.56 86.43 +4.0
Food101 Base 90.10 88.33 88.55 90.70 90.98 90.71 91.00 +0.3
Novel 91.22 82.26 82.35 91.29 91.54 92.05 91.80 -0.9
HM 90.66 85.19 85.36 90.99 91.28 91.38 91.41 +0.1
FGVCAircraft Base 27.19 40.44 40.66 33.41 35.6 37.44 43.27 +5.8
Novel 36.29 22.30 23.1 23.71 28.5 35.61 37.77 +2.0
HM 31.09 28.75 29.46 27.74 31.66 36.50 40.34 +3.7
SUN397 Base 69.36 80.60 80.85 79.74 79.96 80.82 82.77 +2.0
Novel 75.35 65.89 65.91 76.86 77.33 78.70 79.07 +0.3
HM 72.23 72.51 72.62 78.27 78.64 79.75 80.91 +1.2
DTD Base 53.24 79.44 80.56 77.01 78.9 80.36 83.87 +3.5
Novel 59.90 41.18 45.30 56.00 57.9 59.18 63.67 +3.5
HM 56.37 54.24 58 64.85 66.8 68.16 72.20 +4.0
EuroSAT Base 56.48 92.19 92.5 87.49 90.2 94.07 94.50 +0.5
Novel 64.05 54.74 55.65 60.04 63.5 73.23 79.60 +6.4
HM 60.03 68.69 69.49 71.21 74.54 82.35 86.43 +4
UCF101 Base 70.53 84.69 84.10 82.33 82.56 83.00 87.47 +4.5
Novel 77.50 56.05 57.35 73.45 75.56 78.66 79.17 +0.5
HM 73.85 67.46 68.21 77.64 78.92 80.77 83.13 +2.5
3.3. Few-shot Learning Experiments
EnPrompt (Ours) consistently provides improvements on all few-shot settings compared to existing approaches. We note that it is effective for EnPrompt to enhance the few-shot image recognition task through our novel designs. Furthermore, we note that EnPrompt achieves relatively larger gains in minimal data cases, such as for $K=2$ for almost all datasets. This finding suggests that EnPrompt achieves better alignment between the visual and language branches in seen class tasks, even with limited training resources. Moreover, EnPrompt demonstrates significant improvements compared to linear probe CLIP.
Table 5. Cross-dataset benchmark evaluation. EnPrompt (Ours) achieves overall favorable performance.
Source Target
ImageNet Caltech101 OxfordPets StanfordCars Flowers102 Food101 Aircraft SUN397 DTD EuroSAT UCF101 Average
Linear probe CLIP 66.73 92.94 89.07 65.29 71.30 86.11 24.87 62.62 44.56 47.69 66.77 65.12
CoOp 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
CLIP-Adapter 71.40 93.85 89.57 64.66 68.85 85.54 18.53 64.35 41.86 46.43 66.77 64.04
CoCoOp 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
PLOT 70.15 94.60 90.23 65.41 71.97 86.32 22.87 67.22 44.99 46.57 68.32 65.85
MaPLe 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
Ours 71.03 93.93 91.20 65.63 71.73 86.40 25.13 67.67 46.47 48.96 69.73 66.69
3.4. Base-to-Novel Generalization Experiments
In Table 4, EnPrompt (Ours) demonstrates significant improvements on all $11$ datasets. In general, existing approaches CoOp and CoCoOp demonstrate better performance than zero-shot CLIP on base classes. However, when it comes to novel classes, these approaches tend to exhibit normal performance. However, EnPrompt significantly enhances the performance on base classes while also improving the accuracy of zero-shot CLIP on novel classes by $2.69\%$. Overall, EnPrompt provides the best-averaged results of $84.71\%$, $76.90\%$, and $80.64\%$ on the base classes, novel classes, and harmonic mean, respectively.
3.5. Cross Dataset Evaluation
To verify the cross-dataset generalization ability, we train EnPrompt (Ours) on the ImageNet dataset with $1,000$ classes, and test it on the remaining $10$ datasets. As shown in Table 5, EnPrompt shows competitive performance and achieves better generalization in $8/10$ over the CoCoOp. Compared with MaPLe, EnPrompt achieves better performance. This indicates that EnPrompt favors better generalization for a wide range of datasets.
3.6. Domain Generalization Experiments
In Table 6, EnPrompt (Ours) consistently outperforms all existing methods on target datasets, with an overall highest average accuracy of $60.74\%$. EnPrompt achieves average gains of $+1.46\%$ over the CoOp. Compared with MaPLe, EnPrompt shows improved performance in all ImageNet variants datasets. This suggests that EnPrompt is focused on improving generalization capability for domain shifts. Furthermore, EnPrompt provides the highest accuracy of $51.45\%$ on ImageNetA (Hendrycks et al., 2021b).
Table 6. Domain generalization. These approaches are trained on imageNet and tested on datasets with domain shifts.
Source Target
ImageNet -V2 -S -A -R Avg.
Linear probe CLIP 66.73 60.83 46.15 47.77 73.96 57.18
CoOp 71.51 64.2 47.99 49.71 75.21 59.28
CLIP-Adapter 71.40 64.5 47.72 49.75 75.55 59.38
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.91
PLOT 70.15 64.17 49.15 50.83 76.5 60.16
MaPLe 70.72 64.07 49.15 50.9 76.98 60.27
Ours 71.03 64.3 49.5 51.45 77.83 60.77
4. Ablative Analysis
4.1. Comparison of Different Templates
We conducted an experiment using various templates as inputs to the model. Remarkably, we observed that the HM values generated by different templates were very close to each other (see Table 7). The prompt templates can be relatively diversified because of the relatively stronger learning ability of EnLa on unseen tasks, due to the black box (Zhang et al., 2020a) nature of EnLa models.
Table 7. Comparison of different templates on Flowers102.
Templates Base Acc Novel Acc. HM GAP
“a drawing of a” 98.3 77.10 86.42 $\pm$0.05
“a painting of the” 97.2 77.7 86.39 $\pm$0.05
“a photo of a” 98.47 77 86.43 /
4.2. Training Strategy
In Table 8, we evaluate the performance of different EnLa training strategies as an ablation. In our approach, the EnLa is implemented as a neural network, which is essential to learn all potentially valuable features during the training stage in order to achieve effective generalization (Li et al., 2021; Zhang et al., 2020b, 2024). To this end, we focus on training the model for more epochs to learn richer features (row-2), resulting in improved generalization performance (Zhang et al., 2021b; Fang et al., 2024).
Table 8. Comparison of EnLa train strategies. Results are averaged over 11 datasets.
Training epoch Base Acc. Novel Acc. HM
1: EnLa (2 epoch) 83.22 75.91 79.68
2: EnLa (30 epoch) 84.71 76.90 80.64
4.3. Visual Embeddings Length
Figure 4 (left) shows the effect of visual embedding length on the harmonic mean. The visual embeddings are learnable, and the text embeddings are non-learning. We observed a significant decrease in HM when the visual embedding length exceeds $4$. It suggests that too many learnable visual parameters inherently decrease the ability of V-L alignment. Overall, we have determined that using $4$ visual embeddings is the most suitable method.
Figure 4. Ablation study for embedding length and learning depth of image encoder on Flowers102 and UCF101.
4.4. Visual Learning Depth
In Figure 4 (right), we ablate on the visual learning depth for HM. We apply deep prompting on the image encoder. We note that increasing the learning depth generally increases the performance. This suggests that using more layer depth for pre-trained features provides more rich supervision for the features in our approach. The HM value decreases as the number of visual layers increases to $10$ or more. It indicates excessive fine-tuning of the model, causing it to lose CLIP generalization ability. Overall, EnPrompt achieves maximum performance at a depth of $9$.
4.5. Analyzing of Multi-modal Methods
To validate the effect of EnPrompt (Ours), we conducted the base-to-novel task and non-generalization few-shot task using two-modal designs (row-1) and the method completely opposite to EnPrompt (row-2) in Table 9. The method (row-2) is a visual external layer based on frozen visual embeddings with learning textual embeddings. We observed that the method (row-2) has a lower performance in novel classes than the method (row-3), which is due to the learnable text embeddings of method (row-2) being weak in pre-trained generalization compared to the frozen text prompt. Finally, we observed that the method (row-3) and method (row-2) performed better in the novel classes than method (row-1) while still maintaining a non-generalization few-shot task. This is attributed to the fact that method (row-1) has neither learnable visual embeddings nor learnable text embeddings. It indicates that prompt learning is crucial for adapting pre-trained VLMs for generalization.
Table 9. We evaluate EnPrompt with different multi-modal designs. The V-L connection is the default setting.
Multi-modal Methods Base Novel HM Few-shot Task
1: Visual + Text 82.20 75.5 78.71 82.05
2: Opposite Ours (Visual) 84.15 75.20 79.43 82.69
3: Ours (Text) 84.71 76.90 80.64 83.27
4.6. Inference Stage Computational Cost
In Table 10, we show the compute cost analysis of EnPrompt (Ours) and compare it with text embedding learning approaches. Although EnPrompt uses the EnLa, its overall parameters exceed only by $0.52\%$ over CLIP. Compared to MaPLe, EnPrompt has fewer parameters and lower coupling. In terms of inference speed, CoCoOp is significantly slower. In contrast, EnPrompt has no such overhead, obtaining a higher performance with less training time. Although CoOp has a small number of parameters, due to the mismatch of V-L alignment, the training time of 10 epochs for CoOp is similar to ours. EnPrompt is simpler than textual multi-prompt input method (PLOT) in textual design component. Further, EnPrompt provides better convergence as it gets better HM as compared to MaPLe in 10 epochs.
Table 10. The compute cost comparison using SUN397 dataset. Training time for all methods is calculated for 10 epochs on a single A6000 GPU.
Method Params Params % CLIP Train time (min) HM
CoOp 2048 0.002 10.88 71.65
CoCoOp 35360 0.03 39.53 75.83
CLIP-Adapter 0.52M 0.41 8.55 72.23
PLOT 8192 0.008 10.85 76.48
MaPLe 3.55 M 2.85 10.58 79.68
Ours 0.65M 0.52 10.21 80.51
5. Related Work
5.1. Prompt Learning in Vision Language Models
Recently, the strong generalization capability of CLIP (Radford et al., 2021) has made it a foundation for many methods (Fang et al., 2019; Cao et al., 2024b, a) in adapting pre-trained VLMs for downstream tasks (Ding et al., 2023; Luo et al., 2023b; Li et al., 2023b, a). Prompt learning (Liu et al., 2021; Lester et al., 2021; Li and Liang, 2021) is a widely used technique in NLP for learning downstream tasks. The use of text prompts, which are instructions provided to the language branch of a Vision-Language model (VLM), is a common practice to enhance task understanding (Gao et al., 2023; Zhang et al., 2021a; Yang et al., 2024b). CoOp (Zhou et al., 2022c) and CoCoOp (Zhou et al., 2022b) fine-tuning the CLIP model specifically for few-shot image recognition by optimizing a continuous set of token embeddings in the language branch. The image-conditional prompt of CoCoOp significantly contributes to enhancing generalization to unseen classes (Fang et al., 2023, 2022a, 2022b) and mitigating the risk of overfitting to the limited labeled data. Moreover, some approaches (Yao et al., 2023a; Lu et al., 2022; Zhu et al., 2023; Yao et al., 2023b) constrain the proposed learnable prompts to contain the essential general knowledge and prior distribution learning. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022b) ensures that the language model attends to relevant visual information when generating predictions. In addition to single-modal prompt tuning (Jia et al., 2022), some approaches (Khattak et al., 2023; Liu et al., 2023) introduce the multi-modal prompt-tuning designs in CLIP to effectively align V-L representations (Pan et al., 2024; Luo et al., 2024; Di et al., 2024). However, their methods have not fully released the potential generalization ability of CLIP.
5.2. Optimal Transport
Optimal Transport (OT) has recently gained significant attention in various theoretical and application tasks to measure the distance between two probability distributions over metric spaces (Guo et al., 2022; Arjovsky et al., 2017; Tanwisuth et al., 2021; Balaji et al., 2020). For instance, Redko et al.(Redko et al., 2019) tackle the target shift problem by aligning domain distributions using the OT framework. PLOT (Chen et al., 2022) introduces a local text feature with textual multi-prompt input. Ours obtain a global text feature for textual one prompt, instead of a local text feature with textual multi-prompt input.
6. Conclusion
Prompt learning is a promising method for adapting pre-trained vision-language models. However, these methods often struggle to tackle the challenge of generalization on unseen tasks effectively. In this paper, we propose a paradigm called EnPrompt with a novel External Layer (EnLa). Specifically, we propose a textual external layer and learnable visual embeddings for adapting VLMs to downstream tasks. To align the textual and visual features, we propose a novel two-pronged approach. (a) We introduce the optimal transport as the discrepancy metric to align the vision and text modalities. (b) We introduce a novel strengthening feature to enhance the interaction between these two modalities. Extensive experiments clearly demonstrated the effectiveness of our method. In the future, we will extend our effort to more challenging tasks, such as video moment retrieval (Yang et al., 2021, 2022), and investigate the cross-domain generalization ability (Yang et al., 2024a) comprehensively.
7. Acknowledgement
This work was supported in part by National Natural Science Foundation of China (Grant U22A2094 and 62222117), National Key Research and Development Project of China (Grant 2021ZD0110505), and the Zhejiang Provincial Key Research and Development Project (Grant 2023C01043)